OpenStudy (anonymous):

Find all zeros of g(x)=2x^3+x^2-22x+24 given that x=2 is a zero

OpenStudy (anonymous):

Can I use synthetic division here?

OpenStudy (anonymous):

factor as \((x-2)\times (\text{something})\) and then use the quadratic formula to find the zeros of the "something" you can find it by dividing using synthetic division

OpenStudy (anonymous):

yes, use synthetic division is easiest method hardest method is long division medium method is thinking about what it has to be

OpenStudy (anonymous):

Okay. How about this..."Find all factors of x^3-x^2-5x-3 given that (x+1) is a factor"

OpenStudy (anonymous):

Can I use synthetic division there too?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So would the answer be in standard form?

OpenStudy (anonymous):

\[x^3-x^2-5x-3 =(x+1)(whatever)\]

OpenStudy (anonymous):

actually for this one thinking would not be too hard

OpenStudy (anonymous):

\[x^3-x^2-5x-3 =(x+1)(x^2+bx-3)\] you only need to find the \(b\) because the first and last term are obvious

OpenStudy (anonymous):

since \(-3x+bx=-5x\) that tells you \(b=-2\) giving \[x^3-x^2-5x-3 =(x+1)(x^2-2x-3)\]

OpenStudy (anonymous):

but you can use synthetic division if you find it easier

OpenStudy (anonymous):

But it says I have to find the factors..so once I finish with the synthetic division and I have my answer in standard form, do I just factor out whatever I got from the synthetic division?

OpenStudy (anonymous):

So I got this after doing synthetic x^2-2x-3 So would I factor that?

OpenStudy (anonymous):

yeah that one should be easy right?

OpenStudy (anonymous):

Yeah, sorry about my stupidity. I have a test tomorrow and I'm scrambling to learn this pellet

OpenStudy (anonymous):

pellet? I typed in sh**